In the circuit below, determine the value of the voltage across the capacitor, vc, 1 time constant after the sources change (Determine the value of the voltage across the capacitor, Vc, 1 time constant after the sources changes (at t=0).
Answer #1

Civen the ciscuit, 2511) c 10046-t)v a looult) pasu (t) 100V 25v - tro to At t=0 le circuit is in steady state = Copacitor IVoltage across Capacitor, V(t) = 2600)+ [2960) – 260) Jett = 25+ [100-as Jett = 25+ 75ėttoo ie, at, tz 1 time constant = 7. 2if you are facing any difficulty regarding the solution please comment.